OpenStudy (anonymous):

what is the height of a cylinder where volume =48, and the radius is 2?

OpenStudy (shamim):

u know the equation\[v= \pi \times r ^{2} \times h\]

OpenStudy (shamim):

here v= vollume of cyllinder r=radius of cyllinder h= heighr of cyllinder

OpenStudy (shamim):

so 48= 3.14 *2^2*h

OpenStudy (shamim):

solve it

OpenStudy (anonymous):

I got that far but it does not match any of the answers I was given of either 12, 24, 36 or 48. what am I doing wrong?

OpenStudy (shamim):

ok

OpenStudy (shamim):

can u tell me about units given

OpenStudy (anonymous):

That is all that I am given. V=48 and R=2 what is the H. Possible answers - 12, 24, 36 and 48. I know 48=3.14*2*2*h which gives me 48=12.56*h but nothing works out to any of the answers that are possible????